OpenStudy (anonymous):

Solve: 5(x + 8) = 7(x − 2) 17 −17 27 −27 Question 2 (Multiple Choice Worth 3 points) Solve: 2x = −18 −8 9 −6 −9 Question 3 (Multiple Choice Worth 3 points) Solve: −3x − 7 = 17 −3.3 4 8 −8 Question 4 (Multiple Choice Worth 3 points) 5y = 5y is an example of which algebraic property? Symmetric Property Transitive Property Reflexive Property Commutative Property of Multiplication Question 5 (Multiple Choice Worth 3 points) Solve: x − 7 = 14 7 −7 21 −21

OpenStudy (anonymous):

5(x + 8) = 7(x − 2) 5x + 40 = 7x - 14 40-14 = 7x-5x 26 = 2x x = 26/2 x = 13

OpenStudy (anonymous):

tomo: 40+14 = 7x-5x ! q1: x=17 q2: x=-9 q3: x = 8 q4: q5: x=21